# ParallelComp: Parallel Long-Context Compressor for Length Extrapolation

### v1

@@ Appendix A Proof for Parallel Attention Collapse
Consider a parallel attention mechanism where:
  • The input sequence $X\in\mathbb{R}^{N\times d}$ is divided into $C$ chunks, where each chunk $X^{c}\in\mathbb{R}^{w\times d}$ contains at most $w$ tokens, for $c\in[C]$.
  • Define the query, key, and value matrices for each chunk as:
    $Q^{c}=f_{Q}(X^{c}),\quad K^{c}=f_{K}(X^{c}),\quad V^{c}=f_{V}(X^{c}),$
    where $Q^{c},K^{c},V^{c}\in\mathbb{R}^{w\times d}$.
  • The local attention matrix for chunk $c$ is:
    $A^{c}_{\mathfrak{l}}=\operatorname{Softmax}\left(\frac{Q^{c}{K^{c}}^{\top}}{\sqrt{d}}\right)\in\mathbb{R}^{w\times w}.$
  • Let $\mathcal{S}_{\epsilon}^{(c)}$ denote the set of effective entries in the normalized attention matrix row for chunk $c$, where:
    $\mathcal{S}_{\epsilon}^{(c)}(A^{c}_{\mathfrak{l}}[i,:])=\{j\mid A^{c}_{\mathfrak{l}}[i,j]>\epsilon\},\quad i\in[w].$
  • Assume the sparsity parameter $R=O(\sqrt{\log(w)})$.
Then, for any $\epsilon>0$, with high probability ($1-\delta$), the number of ineffective entries in each row satisfies:
@@
In the parallel attention setting, the sparsity of the local attention mechanism is analyzed within each chunk. For a specific chunk $c$, the softmax scaling property governs the behavior of the attention matrix $A^{c}_{\mathfrak{l}}$. The sparsity threshold $\epsilon$ for effective entries in $A^{c}_{\mathfrak{l}}$ can be expressed as:
$\epsilon\geq\exp\left(O(R)\cdot\sqrt{\log(w\cdot(w-k)/\delta)}\right),$
where $k$ represents the number of ineffective entries in a row of the attention matrix. This inequality indicates that as the chunk size $w$ increases, the threshold for retaining effective entries becomes stricter, reducing the number of effective entries.
Rearranging the inequality, we derive an upper bound on $k$, the number of effective entries:
Substituting this bound into the definition of $|\mathcal{S}_{\epsilon}^{(c)}|$, we find that the number of ineffective entries is:
$|\mathcal{S}_{\epsilon}^{(c)}(A^{c}_{\mathfrak{l}}[i,:])|\geq w-k.$
As $w\to\infty$, the remaining effective entries approach $w-1$, as $R=O(\sqrt{\log(w)})$ ensures that the sparsity growth remains bounded. Thus, the number of ineffective entries in each row satisfies:
which completes the proof. ∎
@@
Methods Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code Avg. NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PRe Lcc RB-P Max Length 84123 24204 17727 20325 19001 20520 60515 34477 16271 13049 26756 21884 32699 17158 37628 58822 30657 Llama2-7B-chat-hf(4k) No-eviction 23.20 17.50 37.07 38.67 32.68 20.22 25.00 22.79 25.84 64.00 84.63 40.67 4.00 31.50 59.37 58.53 36.60 Sink-eviction-layer-1-8 23.75 18.69 38.41 39.86 32.91 20.75 24.86 22.10 25.56 63.00 84.42 40.78 4.50 30.00 54.67 59.30 36.47 Sink-eviction-layer-9-16 23.34 19.10 38.21 38.73 30.42 21.04 25.31 21.86 25.16 62.00 85.53 41.26 3.00 29.50 56.75 58.31 36.22 Sink-eviction-layer-17-24 24.46 17.75 36.84 38.79 30.59 19.67 25.42 22.20 25.58 62.00 85.35 40.24 4.00 28.50 58.41 58.17 36.12 Sink-eviction-layer-25-32 23.87 18.40 35.91 38.96 31.02 20.21 25.32 22.00 25.81 64.00 84.19 39.77 3.50 30.00 58.58 58.21 36.23 Recency-eviction-layer-1-8 22.71 16.95 35.24 36.14 30.60 17.19 25.21 22.11 26.22 59.00 68.00 40.03 2.50 31.00 58.07 51.47 33.90 Recency-eviction-layer-9-16 24.95 13.54 35.67 34.13 30.69 17.77 25.14 22.85 25.43 54.50 79.23 39.16 5.00 27.50 57.73 57.57 34.43 Recency-eviction-layer-17-24 21.68 15.17 34.97 32.79 26.93 13.95 25.29 22.10 25.42 62.50 80.47 38.60 6.00 34.00 58.31 57.76 34.75 Recency-eviction-layer-25-32 24.15 17.32 37.82 36.76 29.86 18.48 25.21 22.06 25.67 64.00 83.20 36.67 5.00 30.50 56.13 56.52 35.58 Middle-eviction-layer-1-8 22.41 16.84 37.94 39.99 33.24 19.62 24.74 22.02 25.80 63.50 83.67 40.00 5.00 32.50 59.22 56.71 36.45 Middle-eviction-layer-9-16 22.96 17.63 37.39 40.51 30.68 21.09 25.14 21.94 25.66 62.00 85.02 40.57 4.00 35.50 59.28 57.88 36.70 Middle-eviction-layer-17-24 21.72 17.06 35.88 39.99 31.76 19.06 25.00 22.20 25.77 58.00 83.62 40.02 4.50 38.00 59.38 57.98 36.25 Middle-eviction-layer-25-32 21.64 17.04 36.32 40.51 32.80 19.14 25.07 22.14 25.86 63.50 84.39 40.50 3.00 30.00 59.15 57.86 36.18 All-eviction-layer-1-8 0.33 0.05 1.24 0.32 0.68 0.38 1.76 3.25 1.59 1.50 5.40 1.75 0.41 0.50 23.61 12.17 3.43 All-eviction-layer-9-16 1.60 1.97 5.06 0.65 1.44 0.81 21.75 36.22 1.81 35.00 30.77 10.54 3.00 0.50 33.25 22.89 12.95 All-eviction-layer-17-24 12.20 10.26 16.30 20.30 18.10 8.54 13.63 20.43 17.61 49.00 11.93 29.74 5.50 26.00 41.92 23.67 20.32 All-eviction-layer-25-32 11.19 7.62 11.11 12.49 8.83 1.79 11.45 16.21 12.87 43.00 30.77 9.94 3.50 22.00 24.42 22.19 15.59 Ours-calibration 24.95 19.07 38.16 39.53 32.62 22.64 25.42 22.82 26.01 63.00 85.41 40.36 5.00 32.50 59.04 58.84 37.21
@@ Appendix C Hyperparameter
The Dynamic-PI method interpolates dynamically according to the length of the input token. NTK-Aware refer to (LocalLLaMA, 2023b) and the maximum length is set to 280k. ChunkLlama, InfLLM and AttenCalibration-NTK use hyperparameters from open source repositories. About our method, when performing parallel KV Cache compression, we use last 8 token’s cumulative attention scores to compress the KV cache size within each chunk to 2000. On Longbench, we retain 3 chunks from the priority queue, while on InfiniteBench, we retain 1 chunk for retrieval tasks and 3 chunks for other tasks from the priority queue. In our setup, the total number of chunks processed in parallel and the chunks in the priority queue is 23. In all datasets, the context length of each chunk, including the query, is the maximum pre-training length of the model. $\tau$ is set to 0.1 for llama2 and 0.3 for llama3. $R_{s}$ is obtained from the first 200 tokens of the chunk, $R_{r}$ is obtained from the last 200 tokens of the chunk, and the remaining part of the chunk obtains $R_{m}$. All experiments are performed on 80G A100.

### v2

@@ A.5 Sparsity of Global Attention
For any fixed $\epsilon>0$ and $R=O(\sqrt{\log(Cw+w_{q})})$, as $w_{q}\to\infty$,
$\lim_{w_{q}\to\infty}|\mathcal{S}_{\epsilon}(A_{\mathfrak{g}}[i,:])|=(Cw+w_{q})-k_{\mathfrak{g}},\qquad k_{\mathfrak{g}}=o(Cw+w_{q}).$
@@
Since the global attention decays exponentially in distance, as in the local case the number of entries above $\epsilon$ is $k_{\mathfrak{g}}=o(Cw+w_{q})$ for $R=O(\sqrt{\log(Cw+w_{q})})$, ensuring overall sparsity as sequence length grows. ∎
Discussion.
While global attention has a larger support span, attention values still decay exponentially with distance; only a vanishingly small fraction of entries exceed a fixed threshold for sufficiently large $R$. This structure ensures the computational tractability and sparsity of the attention matrices, where $R$ modulates the effective receptive field between local and global modeling.
@@ Discussion.
Methods Single-Document QA Multi-Document QA Summarization Few-shot Learning Synthetic Code AVG NtrvQA Qasper MF-en HotpotQA 2WikiMQA Musique GovReport QMSum MultiNews TREC TriviaQA SAMSum PCount PRe Lcc RB-P Max Length 84123 24204 17727 20325 19001 20520 60515 34477 16271 13049 26756 21884 32699 17158 37628 58822 30657 Llama2-7B-chat-hf(4k) No-eviction 23.20 17.50 37.07 38.67 32.68 20.22 25.00 22.79 25.84 64.00 84.63 40.67 4.00 31.50 59.37 58.53 36.60 Sink-eviction-layer-1-8 23.75 18.69 38.41 39.86 32.91 20.75 24.86 22.10 25.56 63.00 84.42 40.78 4.50 30.00 54.67 59.30 36.47 Sink-eviction-layer-9-16 23.34 19.10 38.21 38.73 30.42 21.04 25.31 21.86 25.16 62.00 85.53 41.26 3.00 29.50 56.75 58.31 36.22 Sink-eviction-layer-17-24 24.46 17.75 36.84 38.79 30.59 19.67 25.42 22.20 25.58 62.00 85.35 40.24 4.00 28.50 58.41 58.17 36.12 Sink-eviction-layer-25-32 23.87 18.40 35.91 38.96 31.02 20.21 25.32 22.00 25.81 64.00 84.19 39.77 3.50 30.00 58.58 58.21 36.23 Recency-eviction-layer-1-8 22.71 16.95 35.24 36.14 30.60 17.19 25.21 22.11 26.22 59.00 68.00 40.03 2.50 31.00 58.07 51.47 33.90 Recency-eviction-layer-9-16 24.95 13.54 35.67 34.13 30.69 17.77 25.14 22.85 25.43 54.50 79.23 39.16 5.00 27.50 57.73 57.57 34.43 Recency-eviction-layer-17-24 21.68 15.17 34.97 32.79 26.93 13.95 25.29 22.10 25.42 62.50 80.47 38.60 6.00 34.00 58.31 57.76 34.75 Recency-eviction-layer-25-32 24.15 17.32 37.82 36.76 29.86 18.48 25.21 22.06 25.67 64.00 83.20 36.67 5.00 30.50 56.13 56.52 35.58 Middle-eviction-layer-1-8 22.41 16.84 37.94 39.99 33.24 19.62 24.74 22.02 25.80 63.50 83.67 40.00 5.00 32.50 59.22 56.71 36.45 Middle-eviction-layer-9-16 22.96 17.63 37.39 40.51 30.68 21.09 25.14 21.94 25.66 62.00 85.02 40.57 4.00 35.50 59.28 57.88 36.70 Middle-eviction-layer-17-24 21.72 17.06 35.88 39.99 31.76 19.06 25.00 22.20 25.77 58.00 83.62 40.02 4.50 38.00 59.38 57.98 36.25 Middle-eviction-layer-25-32 21.64 17.04 36.32 40.51 32.80 19.14 25.07 22.14 25.86 63.50 84.39 40.50 3.00 30.00 59.15 57.86 36.18 All-eviction-layer-1-8 0.33 0.05 1.24 0.32 0.68 0.38 1.76 3.25 1.59 1.50 5.40 1.75 0.41 0.50 23.61 12.17 3.43 All-eviction-layer-9-16 1.60 1.97 5.06 0.65 1.44 0.81 21.75 36.22 1.81 35.00 30.77 10.54 3.00 0.50 33.25 22.89 12.95 All-eviction-layer-17-24 12.20 10.26 16.30 20.30 18.10 8.54 13.63 20.43 17.61 49.00 11.93 29.74 5.50 26.00 41.92 23.67 20.32 All-eviction-layer-25-32 11.19 7.62 11.11 12.49 8.83 1.79 11.45 16.21 12.87 43.00 30.77 9.94 3.50 22.00 24.42 22.19 15.59 Ours-calibration 24.95 19.07 38.16 39.53 32.62 22.64 25.42 22.82 26.01 63.00 85.41 40.36 5.00 32.50 59.04 58.84 37.21
@@ Appendix C Hyperparameter
The Dynamic-PI method interpolates dynamically according to the length of the input token. NTK-Aware refer to (LocalLLaMA, 2023b) and the maximum length is set to 128. ChunkLlama, InfLLM and AttenCalibration-NTK use hyperparameters from open source repositories. About our method, when performing parallel KV Cache compression, we use last 8 token’s cumulative attention scores to compress the KV cache size within each chunk to 2000/4000 for llama2/llama3. For the hyperparameter $\tau$, on Longbench, we retain 3 chunks from the priority queue except for PRe, in which dataset we retain only 1 chunk. On InfiniteBench, we retain 1 chunk for retrieval tasks and 3 chunks for other tasks from the priority queue. In all datasets, the context length of each chunk, including the query, is the maximum pre-training length of the model. $R_{s}$ is obtained from the first 100 tokens of the chunk, $R_{r}$ is obtained from the last 100 tokens of the chunk, and the remaining part of the chunk obtains $R_{m}$. All experiments are performed on 8 64G AMD Instinct MI210.
Appendix D Comprehensive Evaluation
This section presents additional experiments and implementation details to further contextualize and substantiate the claims in the main text. Specifically, we investigate comparisons with Retrieval-Augmented Generation baselines, provide extended benchmarks with APE (Yang et al., 2025b) and StarAttention (Acharya et al., 2024), report on hyperparameter sensitivity, discuss latency and memory efficiency, and elaborate on our design choices for attention bias sparsification.
D.1 Comparison with Retrieval-Augmented Generation Method
Model QM QASP MSQ HQA MFQA AVG
ChatQA-2 11.64 28.85 27.81 53.81 51.02 34.63
ChatQA-2 w/ RAG 13.20 28.85 29.77 57.81 51.15 36.16
Ours 24.18 39.05 33.25 49.58 42.66 37.74
Table 9: Performance comparison with RAG and non-RAG baselines on LongBench.
Model KV Retrieval Numbe String Passkey En.MC AVG
ChatQA-2 72.00 100.00 100.00 64.19 84.05
ChatQA-2 w/ RAG N/A N/A N/A N/A N/A
Ours 92.80 99.83 100.00 54.59 86.81
Table 10: Performance comparison on InfiniteBench. RAG methods completely fail on InfiniteBench, so we do not provide further results.
To clarify the role of length extrapolation versus retrieval-augmented generation (RAG), we compare the proposed method to leading RAG-enhanced models such as ChatQA-2 (Xu et al., 2024) on representative benchmarks. Table 9 and Table 10 summarize results on LongBench and InfiniteBench, respectively. The proposed method demonstrates strong robustness and competitive or superior performance in challenging long-context retrieval scenarios such as InfiniteBench, where RAG-based methods may encounter instability or diminished effectiveness.
D.2 Comparison with APE and StarAttention
Method NARR QAS MUL HOPT 2WKI MUS GOV QMS NEWS TREC TRIV SSM PCNNT PREN LCC REP AVG
APE 23.63 39.11 50.06 49.47 43.70 25.99 27.78 22.79 11.22 43.50 90.17 9.79 0.50 59.00 23.93 24.28 34.06
StarAttn 3.74 11.90 24.81 14.17 14.37 8.19 34.90 22.54 27.11 65.33 87.84 43.71 3.80 65.17 50.54 45.40 32.72
ParallelComp 29.45 45.98 50.67 48.36 46.56 23.32 32.60 24.29 27.34 38.50 86.72 25.93 0.05 95.00 14.15 21.42 38.15
Table 11: Comparison with APE and StarAttention denoted as StarAttn on LongBench. Temperature and scaling factors for APE are indicated as APE $T+S$. In our experiments, we set APE’s temperature to 0.5 and scaling factor to 0.8; for StarAttention, the chunk size is set to 2K. In ParallelComp, we reuse 6K position encodings to facilitate length extrapolation. All evaluations are conducted on the Llama-3.1-8B-Instruct base model with a KV cache size of 24K.
To comprehensively evaluate our approach in the context of existing chunked long-context processing methods, we conduct extensive experiments comparing ParallelComp with both APE (Yang et al., 2025b) and StarAttention (Acharya et al., 2024). APE leverages a shared prefix to minimize distributional disparities, incorporates a low-temperature mechanism to sharpen attention, and utilizes a scaling factor to compensate for temperature changes. Its objective is to better align the attention patterns between parallel and sequential encoding. StarAttention is designed for chunk-based training of models with long contexts. At each generation step, it recalculates attention for every chunk, whereas ParallelComp computes attention once during the prefill phase and then efficiently reuses the compressed KV cache for subsequent generation. This distinction leads to substantial improvements in computational efficiency.
We first compared the performance of different methods under the same KV cache budget. Hyperparameters are selected according to those reported in the original publications or official releases. Table 11 summarizes the representative results, where all models are assessed using a standardized evaluation infrastructure. Our analysis emphasizes the memory bottleneck encountered during length extrapolation. This experiment emphasizes the memory bottleneck encountered during length extrapolation. The other two methods are forced to truncate the input during extrapolation, resulting in significantly lower performance on certain tasks compared to our approach.
Method NARR QAS MULT HOPT 2WKI MUS GOV QMS NEWS TREC TRIV SSM PCNNT PREN LCC REP AVG APE0.8+0.8 25.92 41.99 53.79 53.64 50.54 26.46 30.15 25.42 20.68 50.50 88.70 9.72 6.50 89.00 16.71 25.78 38.47 APE0.5+0.8 23.63 39.11 50.06 49.47 43.70 25.99 27.78 22.79 11.22 43.50 90.17 9.79 0.50 59.00 23.93 24.28 34.06 APE0.2+0.8 18.83 26.53 41.70 44.63 35.91 17.71 24.31 20.14 7.96 35.75 88.54 9.72 1.50 34.50 23.86 23.22 28.43 APE0.8+0.4 9.04 11.48 19.59 31.41 24.68 10.16 5.32 13.80 8.20 0.50 87.06 9.71 0.00 7.00 13.72 16.51 16.76 APE0.5+0.4 7.59 9.74 16.13 31.89 25.72 9.62 5.20 9.56 8.19 0.00 87.60 9.69 0.00 5.00 14.26 15.58 15.99 APE0.2+0.4 4.90 8.97 13.99 29.71 26.66 8.74 5.16 9.45 8.26 0.50 87.57 9.71 0.00 4.00 14.20 16.24 15.50 StarAttn4K 3.74 11.90 24.81 14.17 14.37 8.19 34.90 22.54 27.11 65.33 87.84 43.71 3.80 65.17 50.54 45.40 32.72 StarAttn6K 4.65 13.63 21.05 14.47 15.57 6.38 34.80 22.67 26.27 66.00 65.54 47.91 8.00 70.00 56.48 45.42 32.43 ParallelComp 29.45 45.98 50.67 48.36 46.56 23.32 32.60 24.29 27.34 38.50 86.72 25.93 0.05 95.00 14.15 21.42 38.15
Table 12: Performance comparison across different methods. APE X+Y indicates temperature = X and scaling factor = Y. StarAttn4K and StarAttn6K represent StarAttention using chunk sizes of 4K and 6K, respectively.
In order to further compare with the full-size context models StarAttention and APE, we conducted ablation experiments as shown in Table 12. We find that even after carefully tuning the hyperparameters of both models, their average performance only surpasses that of our ParallelComp by 0.32%, which further demonstrates the effectiveness of our method.